use pascals triangle to expand the binomical expression (1+2x)^3?


1
Expert's answer
2022-04-13T09:08:28-0400

Let's write first few lines of Pascal's triangle:

"~~~~~~~~~~~~1\\
~~~~~~~~1~~~~~~~1\\
~~~~1~~~~~~2~~~~~~~1\\
1~~~~~~3~~~~~~~3~~~~~~1"

We can see that coefficients for cubic binomial are 1, 3, 3, 1.

"(1+2x)^3=1\cdot1^3+3\cdot1^2\cdot2x+\\
+3\cdot1\cdot(2x)^2+1\cdot(2x)^3=\\
=1+6x+12x^2+8x^3"